# Optimally truncated WKB approximation for the 1D stationary Schrödinger equation in the highly oscillatory regime

### v1

@@ 3 Error analysis
$f_{N,\varepsilon}$ $=\left(S_{0}^{\prime 2}+a\right)+\Bigg{(}\sum_{\begin{subarray}{c}1\leq n+k\leq N\\ 0\leq n,k\leq N\end{subarray}}\varepsilon^{n+k}S_{n}^{\prime}S_{k}^{\prime}+\sum_{n=0}^{N-1}\varepsilon^{n+1}S_{n}^{\prime\prime}\Bigg{)}$ (52)
$\quad+\Bigg{(}\sum_{\begin{subarray}{c}n+k=N+1\\ 0\leq n,k\leq N\end{subarray}}\varepsilon^{N+1}S_{n}^{\prime}S_{k}^{\prime}+\varepsilon^{N+1}S_{N}^{\prime\prime}\Bigg{)}+\sum_{\begin{subarray}{c}n+k>N+1\\ 0\leq n,k\leq N\end{subarray}}\varepsilon^{n+k}S_{n}^{\prime}S_{k}^{\prime}\>.$ (57)
@@
$\sum_{\begin{subarray}{c}1\leq n+k\leq N\\ 0\leq n,k\leq N\end{subarray}}\varepsilon^{n+k}S_{n}^{\prime}S_{k}^{\prime}$ $=\sum_{n=0}^{N-1}\varepsilon^{n+1}\sum_{j=0}^{n+1}S_{j}^{\prime}S_{n+1-j}^{\prime}$ (60)
@@
with $\alpha_{N,\varepsilon}$ from (17). For $N=0$ the last sum drops. The same estimate holds for $\lVert\beta_{N,\varepsilon}\varphi_{N}^{+}\rVert_{L^{\infty}(I)}$. In particular, since the initial values $\varphi_{0}$ and $\varphi_{1}$ are assumed to be uniformly bounded w.r.t. $\varepsilon$, so is $\varphi_{N}^{WKB}$ in $L^{\infty}(I)$.
@@
For $N=0$ the sum in the exponential function drops, and for $N<2$ the double sum drops.
@@ 3.1 Refined error estimate incorporating quadrature errors
For now, let us assume we are given numerical approximations $\widetilde{S}_{n}^{\pm}$, $n\in\mathbb{N}_{0}$, of the functions $S_{n}^{\pm}$, which satisfy $\lVert\widetilde{S}_{n}^{\pm}-S_{n}^{\pm}\rVert_{L^{\infty}(I)}\leq e_{n}$ with positive constants $e_{n}$. We then define the corresponding “perturbed” WKB approximation as
@@
For $N=0$ the sums in the exponential functions drop, and for $N<2$ the double sum drops.
@@
Let us compare the extended error estimate (3.9) with (3.7). The new (additional) second term inside the square brackets in (3.9) is due to the perturbed functions $\widetilde{S}_{n}^{\pm}$ and includes the approximation error bounds $e_{n}$. In particular, the factor $\sum_{n=0}^{N}\varepsilon^{n-1}e_{n}$ is rather unfavorable, as it is of order $\mathcal{O}(\varepsilon^{-1})$, as $\varepsilon\to 0$. We note that the appearance of this $\mathcal{O}(\varepsilon^{-1})$-term in estimate (3.9) is strongly related to the appearance of the $\mathcal{O}(\varepsilon^{-1})$-terms in [3, Theorem 3.1], [4, Theorem 3.2] and [13, Eq. (35)]. There it implied an upper step size limit $h\leq\bar{h}(\varepsilon)=\varepsilon^{\gamma}$ with some $\gamma\in(0,1)$. Similarly, it would require here some $\varepsilon$-dependent upper bound on the quadrature error $e_{0}$ of $\widetilde{S}_{0}^{\pm}$. We specify this observation in the following remark.
@@ 4.1 Computation of the functions $S_{n}$
As the first step, we compute the derivatives $S_{n}^{\prime}$ through (8)-(10) exactly, employing symbolical computation [^3] As an alternative to (8)-(10), in [22] the authors established an almost explicit formula for the derivatives $S_{n}^{\prime}$, depending on $a$ and its derivatives $a^{\prime},\dots,a^{(n)}$. Although not used here, this approach may prove advantageous with regard to the computational time.. Secondly, we employ a highly accurate quadrature for approximating the integral in (14). For this, we use the well-known Clenshaw-Curtis algorithm [9], which we shall briefly explain in the following.
The basic idea of this method is an expansion of the integrand $f$ in terms of Chebyshev polynomials the integrals of which are known. More precisely, one considers a truncated Chebyshev series for the integrand, i.e. $f(l)\approx\sum_{r=0}^{M}a_{r}T_{r}(l)$, $l\in[-1,1]$, where $T_{r}(l)=\cos(r\arccos(l))$, $r\in\mathbb{N}_{0}$, are the Chebyshev polynomials. Here, the spectral coefficients $a_{r}$ are determined with a collocation method on the Chebyshev collocation points $l_{k}=\cos(k\pi/M)$, $k=0,\dots,M$, by solving the $M+1$ equations $f(l_{k})=\sum_{r=0}^{M}a_{r}\cos\left(\frac{rk\pi}{M}\right)$ for the $a_{r}$, $r=0,\dots,M$. Therefore, the spectral coefficients can be computed by the discrete cosine transformation (DCT) of the function $f$ sampled at the collocations points. We note that the DCT is related to the discrete Fourier transform and can be computed efficiently using the fast Fourier transform algorithm after some preprocessing (e.g., see [24, Chap. 8]).
@@
where the coefficients $b_{r}$ are related to the $a_{r}$, see [9] for the detailed formulas. In [8] it was shown that the Clenshaw-Curtis method approximates integrals of analytic functions with spectral accuracy, i.e., the numerical error is decreasing exponentially with the number of modes $M$.
@@
We note that an alternative and efficient way of approximating the functions $S_{n}$ can be realized without the need for symbolical computation of the derivatives $S_{n}^{\prime}$. Indeed, one can instead employ a spectral method to perform the differentiation of the predecessor $S_{n-1}^{\prime}$ in the recursion (10). For instance, by using the $(M+1)\times(M+1)$ Chebyshev differentiation matrices $\mathbf{D}_{M}$ as described in [24, Chap. 6], one can efficiently approximate the derivative of a function at Chebyshev grid points $l_{k}\in[-1,1]$, $k=0,1,\dots,M$. Thus, to approximate the derivative of a function sampled at transformed Chebyshev points $x_{k}\in[\xi,\eta]$, it is necessary to use the scaled matrix $\widetilde{\mathbf{D}}_{M}:=\frac{2}{\eta-\xi}\mathbf{D}_{M}$. Following recurrence relation (8)-(10), we can therefore approximate the derivatives $S_{n}^{\prime}$ sampled at Chebyshev points $x_{k}$ through the following pointwise definition on the grid:
@@
However, note that approximating $S_{n}^{\prime}$ using (79)-(80) can lead to a rapid accumulation of errors, as repeated numerical differentiation is intrinsically unstable. The reason for this behavior lies in the ill-conditioned Chebyshev differentiation matrices $\mathbf{D}_{M}$. It is known that the condition number of these matrices is of order $\mathcal{O}(M^{2})$ (e.g., see [7, 11]). In a finite precision approach this leads to a big loss, which means that roughly two orders of magnitude are lost in accuracy in each application of the recurrence relation to compute the $S_{n}^{\prime}$. Consequently, it is recommendable to employ this approach primarily for small values of $N$.
@@ 4.2 Truncation of the WKB series
Recall that for a fixed $N\geq 0$, Theorem 3.7 guarantees that $\lVert\varphi-\varphi_{N}^{WKB}\rVert_{L^{\infty}(I)}=\mathcal{O}(\varepsilon^{N})$ as $\varepsilon\to 0$, see also Remark 3.8. In practical applications, however, the situation is exactly the opposite, namely, the small parameter $\varepsilon$ is fixed and $N$ can be chosen freely. Note also that just including more terms into the series (6) does not necessarily reduce the error of the WKB approximation, simply since the asymptotic series is typically divergent. Hence the question arises which choice of $N$ will minimize $\lVert\varphi-\varphi_{N}^{WKB}\rVert_{L^{\infty}(I)}$, often referred to as optimal truncation. In this context, we denote with $N_{opt}=N_{opt}(\varepsilon):=\operatorname{argmin}_{N\in\mathbb{N}_{0}}\lVert\varphi-\varphi_{N}^{WKB}\rVert_{L^{\infty}(I)}$ the optimal truncation order. In general, an optimally truncated asymptotic series is sometimes referred to as superasymptotics (e.g., see [6]). The corresponding error of an optimally truncated series is then typically of the form $\sim\exp(-c/\varepsilon)$, as $\varepsilon\to 0$, with some constant $c>0$.
@@
Treating $N$ as a continuous variable for a moment, we find the minimum of
@@
$\lVert\varphi-\varphi_{N}^{WKB}\rVert_{L^{\infty}(I)}\leq\frac{C}{\varepsilon^{2}}\exp\left(-\frac{r}{\varepsilon}\right)\>,$ (86)
@@
$\varepsilon^{N}K_{2}^{N+1}(N+1)^{N+1}\leq\frac{1}{\varepsilon}\left(\frac{c}{\operatorname{e}}\right)^{\left\lfloor\frac{c}{\operatorname{e}K_{2}\varepsilon}\right\rfloor}\leq\frac{\operatorname{e}}{c\varepsilon}\left(\frac{c}{\operatorname{e}}\right)^{\frac{c}{\operatorname{e}K_{2}\varepsilon}}=\frac{\operatorname{e}}{c\varepsilon}\exp\left(-\frac{r}{\varepsilon}\right)\>,$ (90)
with $r:=\frac{c\ln(\operatorname{e}/c)}{\operatorname{e}K_{2}}>0$. Finally, the second term in (89) can be estimated as follows:
@@
$=\frac{c}{(\operatorname{e}-c)K_{2}\varepsilon^{2}}\exp\left(-\frac{r}{\varepsilon}\right)\>.$ (91)
We observe that the r.h.s. of (90) can be bounded by the r.h.s. of (4.2) (up to a multiplicative constant) for $\varepsilon\in(0,\tilde{\varepsilon}_{0}]$. Thus, the claim follows. ∎
@@
$\lVert\varphi-\varphi_{N_{opt}}^{WKB}\rVert_{L^{\infty}(I)}=\mathcal{O}(\varepsilon^{-2}\exp(-r/\varepsilon))\>,\quad\varepsilon\to 0\>,\quad r>0\>.$ (92)
@@
Table 1: Terminology for the different truncation orders mentioned in Section 4. The numbers $\widehat{N}_{opt}$ and $\widehat{N}_{heu}$ are predictions for $N_{opt}$ and ${N}_{heu}$ by means of (3.7) and (82), respectively.
@@ 5 Numerical simulations
In this section we present several numerical simulations to illustrate some of the theoretical results we derived in Section 3. To this end, we will compute the (perturbed) WKB approximation as described in Section 4.1. That is, the functions $S_{n}^{\prime}$ are pre-computed symbolically and are then integrated numerically using the Clenshaw-Curtis algorithm based on a Chebyshev grid with $M+1$ grid points, where $M$ will be specified later. All computations are carried out using Matlab version 9.13.0.2049777 (R2022b). Further, since we are dealing with very small errors for the WKB approximation, especially when investigating the optimal truncation order, we use the Advanpix Multiprecision Computing Toolbox for Matlab [1] with quadruple-precision to avoid rounding errors.
@@ 5.1 Example 1: Airy equation
Here, $\operatorname{Ai}$ and $\operatorname{Bi}$ denote the Airy functions of first and second kind, respectively (e.g., see [21, Chap. 9]). Note that for this example, where $a(x)=x$, the derivatives $S_{n}^{\prime}$ are given by powers of $x$ (up to a constant factor). Hence, the functions $S_{n}^{\pm}$ can be computed exactly from (14); however, we shall use them here only as reference solutions for the approximations $\widetilde{S}_{n}^{\pm}$. Indeed, for a fixed number $M+1$ of Chebyshev grid points, we are then able to compute explicitly the approximation error $\lVert S_{n}^{\pm}-\widetilde{S}_{n}^{\pm}\rVert_{L^{\infty}(I)}=:e_{n}$. Since $\widetilde{S}_{n}^{\pm}$ is only available at the grid points, we actually compute the discrete analog of this norm.
@@
On the left of Figure 2 the real part of $\varphi_{exact}$ is plotted for the choice $\varepsilon=2^{-8}$, which illustrates well the highly oscillatory behavior of the solution.
@@
Figure 2: Left: Real part of the exact solution (94) of IVP (93) for the choice $\varepsilon=2^{-8}$. Right: $L^{\infty}(I)$-norm of $S_{n}^{\prime}$ and $\widetilde{S}_{n}$ as functions of $n$, for the example $a(x)=x$ on the interval $[1,2]$.
@@
Let us first investigate numerically the result from Corollary 3.3. For this, let us compute a constant $K_{2}$, as indicated by the proof of Corollary 3.3 and Remark 3.4. Indeed, by using the minimization strategy from (46), we find that (note that here $S_{0}^{\prime}(z)=\pm\operatorname{i}\sqrt{z}$; $\delta=1$)
$K_{2}=\frac{\operatorname{e}}{2\operatorname{e}-1}\min_{0<\delta\leq 1}\frac{\sqrt{2+\delta}}{\delta}=\frac{\sqrt{3}\operatorname{e}}{2\operatorname{e}-1}\approx 1.0612$ (95)
@@
is a suitable constant within the context of Corollary 3.3. On the right of Figure 2, we present the $L^{\infty}(I)$-norms of the functions $S_{n}^{\prime}$ and the approximations $\widetilde{S}_{n}$ when using $M=25$, along with the theoretical bound (39) on $\lVert S_{n}^{\prime}\rVert_{L^{\infty}(I)}$. We observe that the true norms consistently remain below the theoretical bound. Additionally, we include as a dashed line the theoretical bound (39) when replacing $K_{2}$ and $\lVert S_{0}^{\prime}\rVert_{L^{\infty}(G)}$ by the fitted values $K_{2}^{fitted}=10/37\approx 0.27$ and $0.25$, respectively. We observe very good agreement between the norms $\lVert S_{n}^{\prime}\rVert_{L^{\infty}(I)}$ and the dashed line. This demonstrates well that, in the present example, the norms $\lVert S_{n}^{\prime}\rVert_{L^{\infty}(I)}$ grow as Corollary 3.3 suggests, i.e., $\lVert S_{n}^{\prime}\rVert_{L^{\infty}(I)}\sim CK_{2}^{n}n^{n}$ as $n\to\infty$, for some $C,K_{2}>0$. In general, however, this is not the case. We refer to Appendix A and Section 5.3 for an example, where the functions $S_{n}^{\prime}$ and $S_{n}$ even decay, as $n\to\infty$.

### v2

@@ 3 Error analysis
$f_{N,\varepsilon}$ $=\left(S_{0}^{\prime 2}+a\right)+\Bigg{(}{\color[rgb]{0,0,0}\sum_{n=0}^{N}\sum_{k=\max(0,1-n)}^{N-n}\varepsilon^{n+k}S_{n}^{\prime}S_{k}^{\prime}}+\sum_{n=0}^{N-1}\varepsilon^{n+1}S_{n}^{\prime\prime}\Bigg{)}$
$\quad+\Bigg{(}{\color[rgb]{0,0,0}\sum_{n=1}^{N}\varepsilon^{N+1}S_{N+1-n}^{\prime}S_{n}^{\prime}}+\varepsilon^{N+1}S_{N}^{\prime\prime}\Bigg{)}+{\color[rgb]{0,0,0}\sum_{n=2}^{N}\sum_{k=N+2-n}^{N}\varepsilon^{n+k}S_{n}^{\prime}S_{k}^{\prime}}\>.$ (45)
@@
${\color[rgb]{0,0,0}\sum_{n=0}^{N}\sum_{k=\max(0,1-n)}^{N-n}\varepsilon^{n+k}S_{n}^{\prime}S_{k}^{\prime}}$ $=\sum_{n=0}^{N-1}\varepsilon^{n+1}\sum_{j=0}^{n+1}S_{j}^{\prime}S_{n+1-j}^{\prime}$
@@
with $\alpha_{N,\varepsilon}$ from (17). For $N=0$ the last sum is dropped. The same estimate holds for $\lVert\beta_{N,\varepsilon}\varphi_{N}^{+}\rVert_{L^{\infty}(I)}$. In particular, since the initial values $\varphi_{0}$ and $\varphi_{1}$ are assumed to be uniformly bounded w.r.t. $\varepsilon$, so is $\varphi_{N}^{WKB}$ in $L^{\infty}(I)$.
@@
For $N=0$ the sum in the exponential function is dropped, and for $N<2$ the double sum is dropped.
@@ 3.1 Refined error estimate incorporating quadrature errors
For now, let us assume we are given numerical approximations $\widetilde{S}_{n}^{\pm}$, $n\in\mathbb{N}_{0}$, of the functions $S_{n}^{\pm}$ that satisfy $\lVert\widetilde{S}_{n}^{\pm}-S_{n}^{\pm}\rVert_{L^{\infty}(I)}\leq e_{n}$ with positive constants $e_{n}$. We then define the corresponding “perturbed” WKB approximation as
@@
For $N=0$ the sums in the exponential functions drop, and for $N<2$ the double sum is dropped.
@@
Let us compare the extended error estimate (3.9) with (3.7). The new (additional) second term inside the square brackets in (3.9) is due to the perturbed functions $\widetilde{S}_{n}^{\pm}$ and includes the approximation error bounds $e_{n}$. In particular, the factor $\sum_{n=0}^{N}\varepsilon^{n-1}e_{n}$ is rather unfavorable, as it is of order $\mathcal{O}(\varepsilon^{-1})$, as $\varepsilon\to 0$. We note that the appearance of this $\mathcal{O}(\varepsilon^{-1})$-term in estimate (3.9) is strongly related to the appearance of the $\mathcal{O}(\varepsilon^{-1})$-terms in [3, Theorem 3.1], [4, Theorem 3.2] and [16, Eq. (35)]. There it implied an upper step size limit $h\leq\bar{h}(\varepsilon)=\varepsilon^{\gamma}$ with some $\gamma\in(0,1)$. Similarly, it would require here some $\varepsilon$-dependent upper bound on the quadrature error $e_{0}$ of $\widetilde{S}_{0}^{\pm}$ in order to compensate at least the $\mathcal{O}(\varepsilon^{-1})$-error term. In practice this will necessitate to use some finer grid for computing $\widetilde{S}_{0}^{\pm}$, as $\varepsilon$ decreases. We specify this observation in the following remark.
@@ 4.1 Computation of the functions $S_{n}$
As the first step, we compute the derivatives $S_{n}^{\prime}$ through (8)-(10) exactly, employing symbolic computations [^3] As an alternative to (8)-(10), in [25] the authors established an almost explicit formula for the derivatives $S_{n}^{\prime}$, depending on $a$ and its derivatives $a^{\prime},\dots,a^{(n)}$. Although not used here, this approach may prove advantageous with regard to the computational time.. Secondly, we employ a highly accurate quadrature for approximating the integral in (14). For this, we use the well-known Clenshaw-Curtis algorithm [12], which we shall briefly explain in the following.
The basic idea of Clenshaw-Curtis quadrature is to expand the integrand $f$ in terms of Chebyshev polynomials, the integrals of which are known. More precisely, one considers a truncated Chebyshev series for the integrand, i.e., $f(l)\approx\sum_{r=0}^{M}a_{r}T_{r}(l)$, $l\in[-1,1]$, where $T_{r}(l)=\cos(r\arccos(l))$, $r\in\mathbb{N}_{0}$, are the Chebyshev polynomials. Here, the spectral coefficients $a_{r}$ are determined with a collocation method at the Chebyshev collocation points $l_{k}=\cos(k\pi/M)$, $k=0,\dots,M$, by solving the $M+1$ equations $f(l_{k})=\sum_{r=0}^{M}a_{r}\cos\left(\frac{rk\pi}{M}\right)$ for the $a_{r}$, $r=0,\dots,M$. Therefore, the spectral coefficients can be computed by the discrete cosine transformation (DCT) of the function $f$ sampled at the collocations points. We note that the DCT is related to the discrete Fourier transform and can be computed efficiently using the fast Fourier transform algorithm after some preprocessing (e.g., see [27, Chap. 8]).
@@
where the coefficients $b_{r}$ are related to the $a_{r}$, see [12] for the detailed formulas. In [11] it was shown that the Clenshaw-Curtis method approximates integrals of analytic functions with spectral accuracy, i.e., the numerical error decreases exponentially with the number of modes $M$.
@@
We note that an alternative and efficient way of approximating the functions $S_{n}$ can be realized without the need for symbolic computation of the derivatives $S_{n}^{\prime}$. Indeed, one can instead employ a spectral method to perform the differentiation of the predecessor $S_{n-1}^{\prime}$ in the recursion (10). For instance, by using the $(M+1)\times(M+1)$ Chebyshev differentiation matrices $\mathbf{D}_{M}$ as described in [27, Chap. 6], one can efficiently approximate the derivative of a function at Chebyshev grid points $l_{k}\in[-1,1]$, $k=0,1,\dots,M$. Thus, to approximate the derivative of a function sampled at transformed Chebyshev points $x_{k}\in[\xi,\eta]$, it is necessary to use the scaled matrix $\widetilde{\mathbf{D}}_{M}:=\frac{2}{\eta-\xi}\mathbf{D}_{M}$. Following recurrence relation (8)-(10), we can therefore approximate the derivatives $S_{n}^{\prime}$ sampled at Chebyshev points $x_{k}$ through the following pointwise definition on the grid:
@@
However, note that approximating $S_{n}^{\prime}$ using (64)-(65) can lead to a rapid accumulation of errors, as repeated numerical differentiation is intrinsically unstable. The reason for this behavior lies in the ill-conditioned Chebyshev differentiation matrices $\mathbf{D}_{M}$. It is known that the condition number of these matrices is of order $\mathcal{O}(M^{2})$ (e.g., see [10, 14]). In a finite precision approach this leads to a big loss, which means that in each application of the recurrence relation we lose a finite amount of accuracy in the computation of the $S_{n}^{\prime}$ (see Figure 2 for two examples). Consequently, it can be recommended to employ this approach only for small values of $N$.
Figure 2: $L^{\infty}(I)$-norm of the error of the approximation $\bar{S}_{n}^{\prime}$ for the examples $a(x)=x$ and $a(x)=x^{2}$ on the interval $I=[1,2]$. Here, we set $M=20$.
@@ 4.2 Truncation of the WKB series
Recall that for a fixed $N\geq 0$, Theorem 3.7 guarantees that $\lVert\varphi-\varphi_{N}^{WKB}\rVert_{L^{\infty}(I)}=\mathcal{O}(\varepsilon^{N})$ as $\varepsilon\to 0$, see also Remark 3.8. In practical applications, however, the situation is exactly the opposite, namely, the small parameter $\varepsilon$ is fixed and $N$ can be chosen freely. Note also that just including more terms into the series (6) does not necessarily reduce the error of the WKB approximation, simply since the asymptotic series is typically divergent. Hence the question arises which choice of $N$ will minimize $\lVert\varphi-\varphi_{N}^{WKB}\rVert_{L^{\infty}(I)}$, often referred to as optimal truncation. In this context, we denote by $N_{opt}=N_{opt}(\varepsilon):=\operatorname{argmin}_{N\in\mathbb{N}_{0}}\lVert\varphi-\varphi_{N}^{WKB}\rVert_{L^{\infty}(I)}$ the optimal truncation order. In general, an optimally truncated asymptotic series is sometimes referred to as superasymptotics (e.g., see [7]). The corresponding error of an optimally truncated series is then typically of the form $\sim\exp(-c/\varepsilon)$, as $\varepsilon\to 0$, with some constant $c>0$.
@@
Treating $N$ as a continuous variable for the moment, we find the minimum of
@@
$\lVert\varphi-\varphi_{N}^{WKB}\rVert_{L^{\infty}(I)}\leq{\color[rgb]{0,0,0}C\exp\left(-\frac{r}{\varepsilon}\right)}\>,$ (71)
@@
$\varepsilon^{N}K_{2}^{N+1}(N+1)^{N+1}\leq\frac{1}{\varepsilon}\left(\frac{c}{\operatorname{e}}\right)^{\left\lfloor\frac{c}{\operatorname{e}K_{2}\varepsilon}\right\rfloor}\leq\frac{\operatorname{e}}{c\varepsilon}\left(\frac{c}{\operatorname{e}}\right)^{\frac{c}{\operatorname{e}K_{2}\varepsilon}}=\frac{\operatorname{e}}{c\varepsilon}\exp\left(-\frac{{\color[rgb]{0,0,0}\tilde{r}}}{\varepsilon}\right)\>,$ (75)
with ${\color[rgb]{0,0,0}\tilde{r}}:=\frac{c\ln(\operatorname{e}/c)}{\operatorname{e}K_{2}}>0$. Finally, the second term in (74) can be estimated as follows:
@@
$=\frac{c}{(\operatorname{e}-c)K_{2}\varepsilon^{2}}\exp\left(-\frac{{\color[rgb]{0,0,0}\tilde{r}}}{\varepsilon}\right)\>.$ (76)
We observe that the r.h.s. of (75) can be bounded by the r.h.s. of (4.2) (up to a multiplicative constant) for $\varepsilon\in(0,\tilde{\varepsilon}_{0}]$. Thus, the claim follows with $r:=\tilde{r}/2$ and adapting $C$. ∎
@@
$\lVert\varphi-\varphi_{N_{opt}}^{WKB}\rVert_{L^{\infty}(I)}={\color[rgb]{0,0,0}\mathcal{O}(\exp(-r/\varepsilon))}\>,\quad\varepsilon\to 0\>,\quad r>0\>.$ (77)
@@
Table 1: Terminology for the different truncation orders mentioned in Section 4. The numbers $\widehat{N}_{opt}$ and $\widehat{N}_{heu}$ are predictions for $N_{opt}$ and ${N}_{heu}$ by means of (3.7) and (67), respectively. (For numerical values in two concrete examples see Figures 6 and 9.)
@@ 5 Numerical simulations
In this section we present several numerical simulations to illustrate some of the theoretical results we derived in Section 3. To this end, we will compute the (perturbed) WKB approximation as described in Section 4.1. That is, the functions $S_{n}^{\prime}$ are pre-computed symbolically and are then integrated numerically using the Clenshaw-Curtis algorithm based on a Chebyshev grid with $M+1$ grid points, where $M$ will be specified later. All computations are carried out using Matlab version 9.13.0.2049777 (R2022b). Further, since we are dealing with very small errors for the WKB approximation, especially when investigating the optimal truncation order, we use the Advanpix Multiprecision Computing Toolbox for Matlab [1] with quadruple-precision to avoid roundoff errors.
@@ 5.1 Example 1: Airy equation
Here, $\operatorname{Ai}$ and $\operatorname{Bi}$ denote the Airy functions of first and second kind, respectively (e.g., see [24, Chap. 9]). Note that for this example, where $a(x)=x$, the derivatives $S_{n}^{\prime}$ are given by powers of $x$ (up to a constant factor). Hence, the functions $S_{n}^{\pm}$ can be computed exactly from (14); however, we shall use them here only as reference solutions for the approximations $\widetilde{S}_{n}^{\pm}$. Indeed, for a fixed number $M+1$ of Chebyshev grid points, we are then able to compute explicitly the approximation error $\lVert S_{n}^{\pm}-\widetilde{S}_{n}^{\pm}\rVert_{L^{\infty}(I)}=:e_{n}$. Since $\widetilde{S}_{n}^{\pm}$ is only available at the grid points, we actually compute the discrete analog of this norm.
@@
The left panel of Figure 3 shows the real part of $\varphi_{exact}$ for the choice $\varepsilon=2^{-8}$, which illustrates well the highly oscillatory behavior of the solution.
@@
Figure 3: Left: Real part of the exact solution (79) of IVP (78) for the choice $\varepsilon=2^{-8}$. Right: $L^{\infty}(I)$-norm of $S_{n}^{\prime}$ and $\widetilde{S}_{n}$ as functions of $n$ for the example $a(x)=x$ on the interval ${\color[rgb]{0,0,0}I=}[1,2]$.
@@
Let us first investigate numerically the result from Corollary 3.3. For this, let us compute a constant $K_{2}$, as indicated by the proof of Corollary 3.3 and Remark 3.4. Indeed, by using the minimization strategy from (41), we find that (note that here $S_{0}^{\prime}(z)=\pm\operatorname{i}\sqrt{z}$; $\delta_{opt}=\frac{\sqrt{97}-7}{6}\approx 0.4748$)
$K_{2}=\frac{\operatorname{e}}{2\operatorname{e}-1}\min_{0<\delta\leq 1}{\color[rgb]{0,0,0}\frac{\sqrt{2+\delta}}{\delta(1-\delta)}=\frac{\operatorname{e}}{2\operatorname{e}-1}\frac{\sqrt{2+\delta_{opt}}}{\delta_{opt}(1-\delta_{opt})}\approx 3.8653}$ (80)
@@
is a suitable constant within the context of Corollary 3.3. In the right panel of Figure 3, we present the $L^{\infty}(I)$-norms of the functions $S_{n}^{\prime}$ and the approximations $\widetilde{S}_{n}$ when using $M=25$, along with the theoretical bound (39) on $\lVert S_{n}^{\prime}\rVert_{L^{\infty}(I)}$. We observe that the true norms consistently remain below the theoretical bound. Additionally, we include as a dashed line the theoretical bound (39) when replacing $K_{2}$ and $\lVert S_{0}^{\prime}\rVert_{L^{\infty}(G)}$ by the experimentally fitted values $K_{2}^{fitted}=10/37\approx 0.27$ and $0.25$, respectively. [^4] Dividing (39) by $n^{n}$ and taking the logarithm we used a linear approximation to obtain $K_{2}^{fitted}$. We observe very good agreement between the norms $\lVert S_{n}^{\prime}\rVert_{L^{\infty}(I)}$ and the dashed line. This demonstrates well that, in the present example, the norms $\lVert S_{n}^{\prime}\rVert_{L^{\infty}(I)}$ grow as Corollary 3.3 suggests, i.e., $\lVert S_{n}^{\prime}\rVert_{L^{\infty}(I)}\sim CK_{2}^{n}n^{n}$ as $n\to\infty$, for some appropriate constants $C,K_{2}>0$. In general, however, this may not be the case. We refer to Appendix A and Section 5.3 for an example, where the functions $S_{n}^{\prime}$ and $S_{n}$ even decay as $n\to\infty$.